Bank of Springfield Assets Liabilities Reserves $19,800 Deposits $180,000 Loans 160,200 Refer to Table 29-7. If the Bank of Springfield has lent out all the money it can given its level of deposits, then what is the reserve requirement?
Answer #1

Let the reserve requirement be r %.
Reserves = $19,800
Deposits = $180,000

so, r % of deposits = reserves
i.e., (r/100)*180,000 = 19,800
So, r = (19,800*100)/180,000 = 11%

So, the reserve requirement is 11%